OpenStudy (anonymous):

simplify the following expression: 6z^2+25z+25 / 9z^+9z-10

OpenStudy (anonymous):

\[\frac{ 6z^2 +25z+25 }{ 9z^2 +9z-10 }\]

OpenStudy (anonymous):

factor and cancel for this one

OpenStudy (anonymous):

\[\frac{ 6z^2 +25z+25 }{ 9z^2 +9z-10 }\] \[=\frac{(2z+5)(3z+5)}{(3z-2)(3z+5)}\]

OpenStudy (anonymous):

3z-2 / 2z+5

OpenStudy (anonymous):

no, cancel the common factor i wrote above